### v2

# BayesFLo: Bayesian fault localization of complex software systems

Abstract
Software testing is essential for the reliable development of complex software systems. A key step in software testing is fault localization, which uses test data to pinpoint failure-inducing combinations for further diagnosis. Existing fault localization methods have two key limitations: they (i) do not incorporate domain and/or structural knowledge from test engineers, and (ii) do not provide a probabilistic assessment of risk for potential root causes. Such methods can thus fail to confidently whittle down the combinatorial number of potential root causes in complex systems, resulting in prohibitively high testing costs. To address this, we propose a novel Bayesian fault localization framework called BayesFLo, which leverages a flexible Bayesian model for identifying potential root causes with probabilistic uncertainty. Using a carefully-specified prior on root cause probabilities, BayesFLo permits the integration of domain and structural knowledge via the principles of combination hierarchy and heredity, which capture the expected structure of failure-inducing combinations. We then develop new algorithms for efficient computation of posterior root cause probabilities, leveraging recent tools from integer programming and graph representations. Finally, we demonstrate the effectiveness of BayesFLo over existing methods in two fault localization case studies on the Traffic Alert and Collision Avoidance System and the JMP Easy DOE platform.
Keywords: Bayesian modeling, Combinatorial testing, Fault localization, Software testing
1 Introduction
Software testing – the process of executing a program with the intent of finding errors (Myers et al.,, 2004) – is an essential step in the development of robust software applications. Such testing aims to reveal and subsequently fix as many bugs (i.e., faults) as possible prior to the deployment of a software application, thus greatly reducing the likelihood of encountering failures for the end-user. This is crucial in an era where nearly all facets of daily life involve human interaction with software applications. There are, however, two fundamental challenges. First, each software test can be time-consuming to perform. This involves not only running the software system itself, which can be intensive for modern scientific applications (Ji et al., 2024a, ; Ji et al., 2024b, ), but also determining whether the software deviates from its expected behavior. The latter, known as the “oracle problem” (Barr et al.,, 2014), typically requires an independent assessment of numerical accuracy of software outputs (Lekivetz and Morgan,, 2021), which can be very costly. Second, the number of test cases required for thorough software examination can easily be overwhelming. As “bugs [tend to] lurk in corners and congregate at boundaries” (Beizer,, 2003), software testing typically focuses on boundary values and the combinations of inputs, which can grow rapidly. For practical applications, it is thus wholly infeasible to exhaustively test all input combinations (Kumar,, 2019), which can easily require billions of test cases.
These two fundamental challenges open up a world of exciting new statistical directions for this application of rising importance. Such directions can roughly be categorized into two topics. The first is a careful design of test cases to perform with the joint goals of identifying failure settings and diagnosing its underlying root causes. Statistically, this can be viewed as an experimental design problem for software fault diagnosis. There has been notable work on this front. Early work on one-factor-at-a-time designs (Frey et al.,, 2003) explored the problem of unit testing (Runeson,, 2006; Khorikov,, 2020), which focuses on testing small functional code units, e.g., a single function or chunk of code. However, such testing cannot identify bugs that congregate at the interaction of such units, which often arise in large software applications. Another approach is pairwise testing (Bach and Schroeder,, 2004), which examines all pairwise combinations of inputs; test case generation for this setting has been explored in Tai and Lei, (2002). A more general approach is combinatorial testing (Nie and Leung, 2011b, ), which investigates combinations involving more than two inputs. For combinatorial testing, the design of choice is a covering array (CA; Colbourn,, 2004; Dalal and Mallows,, 1998); such designs aim to represent (or “cover”) each combination of inputs (up to a specified order) at least once. CAs are thus ideal for detecting failures from limited test runs; we will discuss CAs in greater detail in Section 2.
With the initial test cases conducted and failures detected, the second direction is fault localization (Wong et al.,, 2023): the use of this test data to pinpoint root causes. This is a highly challenging problem due to the overwhelming number of scenarios to consider for potential root causes. To see why, consider a software application with $I=10$ input factors each with two levels. For this system, there are a total of $\sum_{i=1}^{10}\binom{10}{i}2^{i}=59,048$ input combinations that might be potential root causes; these consist of all single factors, two-factor combinations, and so on. Since each combination is either a root cause or not, this results in a whopping $2^{59048}$ different scenarios to consider for potential root causes! Fault localization requires gauging which of these many scenarios is likely given test set outcomes, which is clearly a computationally intensive task (Wong et al.,, 2023), even for systems with a moderate number of inputs $I$ and few failed test cases. Furthermore, the debugging cost for investigating each such scenario can be considerably high (Boehm,, 1984; Humphrey,, 1995). For example, for critical flight software systems at the Jet Propulsion Laboratory (Bush,, 1990), each debugging activity can cost over thousands of dollars to identify and fix faults.
Due to this combinatorial explosion of potential root causes and high debugging costs, there is thus an urgent need for fault localization methods that, based on initial test case data, can identify a small number of suspicious input combinations for further investigation. On this front, Nie and Leung, 2011a proposed a minimal failure-causing schema and used it to narrow down the search range for potential root causes and to guide subsequent test case generation. Niu et al., (2013) proposed a notion of tuple relationship tree for visualizing the relationships among all input combinations. Such a tree is utilized to eliminate “healthy” combinations and to propose subsequent test cases for further system examination. More recently, Ghandehari, (2016) and Ghandehari et al., (2018) introduced a two-phase approach for finding faulty statements in a software system. Such approaches have been integrated, either in full or in part, within the Covering Array Analysis module in the statistical software package JMP (henceforce called JMP; Jones and Sall,, 2011).
The above body of work, however, has two key limitations that hamper its effectiveness in practical applications. First, they are not probabilistic in nature, and thus offer little insight on the probability of a combination being a root cause given test outcomes. Such probabilities are needed for confident fault localization; they provide a principled measure of risk on the “suspiciousness” of a root cause. A probabilistic approach can further accelerate fault localization by disentangling the many tied combinations given limited test runs; more on this later. Second, existing methods do not permit the integration of domain and/or structural knowledge from test engineers within a probabilistic framework. Here, domain knowledge refers to prior information that may raise or reduce the suspiciousness of certain input factors; this may come, e.g., from historical data, expert judgment or specific characteristics of the system. Structural knowledge, in contrast, captures expected fault behavior from an understanding of the software’s architecture and how components interact (Lekivetz and Morgan,, 2018, 2021). Existing methods that do not integrate such knowledge may fail to whittle down the combinatorial number of potential root causes, which leads to costly and/or delayed fault diagnosis, as we shall see in our applications later.
To address these challenges, we propose a new Bayesian Fault LOcalization (BayesFLo) framework. The main workhorse of BayesFLo is a new Bayesian model on root cause indicators over all possible input combinations, which facilitates easy integration of domain knowledge from test engineers. This model further embeds the desirable principles of combination hierarchy and heredity (Lekivetz and Morgan,, 2021), which capture structural knowledge on software root causes. We show that the integration of such principles, derived from the well-known principles of effect hierarchy and heredity (Wu and Hamada,, 2009) for analyzing experimental data, can accelerate the identification of root causes from limited test cases. A critical challenge for Bayesian computation is the sheer number of considered combinations, which renders the computation of posterior root cause probabilities to be infeasible without careful manipulation. We thus develop a new algorithm for efficient computation of such posterior probabilities, leveraging recent tools from integer programming and graph representations. Finally, we demonstrate the practical advantages of BayesFLo over existing methods in two fault localization case studies, the first for the Traffic alert and Collision Avoidance System (TCAS; Bradley,, 1992) and the other for the Easy DOE platform in JMP (Jones and Sall,, 2011).
The paper is organized as follows. Section 2 outlines our motivating application on fault localization of TCAS, as well as limitations of the state-of-the-art. Section 3 presents the BayesFLo model, and describes the combination hierarchy and heredity principles embedded in its prior specification. Section 4 proposes a novel algorithm for efficient computation of posterior root cause probabilities for potential root cause combinations. Section 5 explores the application of BayesFLo in two practical case studies on the fault localization of TCAS and the JMP Easy DOE platform. Section 6 concludes the paper.
2 Motivating Application: Fault Localization of TCAS
2.1 Background & Challenges
Our motivating application is in the fault localization of the Traffic alert and Collision Avoidance System (TCAS; Harman,, 1989; Williamson and Spencer,, 1989), which is jointly developed by the Federal Aviation Administration (FAA) and the aviation industry. TCAS is an aircraft collision avoidance system, designed to prevent potential mid-air collisions between aircraft (FAA,, 2011). Figure 1 visualizes its operation. TCAS monitors the airspace around an aircraft, and determines a protection volume around the aircraft based on its speed and trajectory. When two aircraft come too close either horizontally or vertically, TCAS alerts the pilots by issuing a resolution advisory. TCAS is currently mandated for U.S. commercial aircraft, and is also installed on many helicopters and turbine-powered general aviation aircraft (FAA,, 2011). Given the catastrophic implications of aircraft collisions, e.g., from recent events (Davies and Yousif,, 2025), it is paramount to mitigate such risks by ensuring the reliability of TCAS systems via comprehensive testing under different scenarios.
Figure 1: Visualizing the protection volume module in the TCAS software system (FAA,, 2011).
Factor Description Levels
1 Cur_Vertical_Sep 299, 300, 601
2 High_Confidence 0, 1
3 Two_of_Three_Reports_Valid 0, 1
4 Own_Tracked_Alt 1, 2
5 Own_Tracked_Alt_Rate 600, 601
6 Other_Tracked_Alt 1, 2
7 Alt_Layer_Value 0, 1, 2, 3
8 Up_Separation 0, 300, 399
9 Down_Separation 640, 740, 840
10 Other_RAC 0, 1, 2
11 Other_Capability 1, 2
12 Climb_Inhibit 0, 1
Table 1: Input factors and their corresponding levels for our motivating TCAS case study.
Testing the TCAS software system has attracted much attention within the software engineering community, and there has been fruitful work on designing test suites for detecting defects in the system (Griesmayer et al.,, 2007; Ghandehari et al.,, 2013). In what follows, we consider a specific TCAS program from the TCAS software system – the Siemens test suite in Griesmayer et al., (2007) – for fault localization. This program takes in $I=12$ input factors (summarized in Table 1 with its corresponding levels), with each factor parameterizing certain characteristics of the aircraft system. Of these $I=12$ factors, nine pertain to the primary aircraft, and the remaining three (with prefix Other in Table 1) are for an incoming aircraft that has entered its protection volume (Xie et al.,, 2013). The latter consists of inputs 6, 10 and 11, which reflect the altitude, resolution advisory and TCAS capability of the incoming aircraft. The former consists of remaining inputs that reflect the operating conditions of the primary aircraft, e.g., its current altitude, altitude rate-of-change, and separation altitudes with the incoming aircraft.
A key challenge here is the expensive testing cost for each run of TCAS. Recall that a “failure” in a software program is when its actual behavior deviate from its expected behavior. Here, the assessment of the latter requires the construction of eleven data collection sites over the U.S. (Bradley,, 1992) and the collection of ground radar data (spanning hundreds of hours) for generating encounter databases. Logic test simulations are performed on thousands of such encounters (Bradley,, 1992), and physical experiments are then conducted to validate its real-world performance. Coupled with the many potential root cause combinations from the $I=12$ input factors, such testing costs make the exhaustive testing of TCAS over all combinations wholly impractical.
An effective solution is to first construct a small designed test set of input settings. An appealing design employed in the literature is a covering array (CA; Colbourn,, 2004; Lekivetz and Morgan,, 2021), which covers each input combination (up to a certain order) at least once. Formally, take an array with dimensions $M\times I$, and suppose its $i$-th column takes $J_{i}$ distinct levels for some integer $J_{i}\geq 2$. Then this array is a CA of strength $s$, if for any choice of $s$ columns, each level combination involving such columns occurs at least once. For software testing, this CA can be used to design a test set with $M$ runs and $I$ inputs, where the $i$-th input has $J_{i}$ levels; the levels in its $m$-th row specify the input settings for the $m$-th test run. Table 2 shows a strength-2 CA for our TCAS case study. Such a design achieves the desired coverage condition with $M=17$ runs, which greatly reduces the number of validation runs on the expensive software system. With this CA, all single inputs and two-factor input combinations are investigated in at least one test run; if there is a bug within such combinations, we would observe a corresponding failure in a test run.
The next step is to use this test run data to identify potential root causes, i.e., the problem of fault localization. The last column in Table 2 shows the outcomes (0 for a passed case, 1 for a failed case) for each test run. A key challenge here is that there are more than $\sum_{i=1}^{12}\binom{12}{i}2^{i}>500,000$ input combinations (e.g., the two-factor combination Cur_Vertical_Sep = 299 and High_Confidence = 0) that may be failure-inducing root causes. Since each combination is either a root cause or not, there are thus a whopping $2^{500,000}$ different root cause scenarios to consider for fault localization! The goal is to parse through these many scenarios to find which are likely given the observed test results. With the costly nature of debugging, this set of likely scenarios should further consist of only a few combinations, to facilitate cost-efficient downstream diagnosis. Another challenge is the need for assessing confidence that an identified suspicious combination is indeed a root cause. This provides test engineers a principled way to decide which combinations are likely root causes that need to be further investigated, and which are likely not root causes and can be safely ignored. Such uncertainty quantification is critical for trustworthy fault diagnosis (Zhou et al.,, 2023). We inspect next existing methods through the lens of these two challenges.
Factor
Run Number 1 2 3 4 5 6 7 8 9 10 11 12 Outcome
1 601 1 1 1 601 1 3 300 640 1 1 0 0
2 300 1 1 1 600 1 0 0 740 0 1 1 0
3 299 0 0 2 600 1 2 300 640 0 1 0 0
4 300 0 0 2 600 2 1 300 740 2 1 0 0
5 601 0 0 2 601 1 3 399 840 1 1 1 0
6 601 1 0 1 601 2 0 300 840 0 2 0 0
7 300 1 0 2 600 2 3 0 840 0 2 0 0
8 299 0 1 2 600 1 3 0 640 2 1 1 0
9 299 1 0 1 601 2 0 399 640 2 1 0 0
10 601 1 1 1 601 1 1 0 740 2 2 1 0
11 299 0 0 2 600 1 3 399 740 0 2 1 0
12 300 1 1 2 601 1 2 399 840 2 2 1 0
13 299 0 1 1 601 1 1 399 640 0 1 1 1
14 299 0 1 1 600 1 1 300 840 1 1 1 0
15 300 0 0 2 600 1 0 399 640 1 2 1 1
16 601 1 1 1 600 2 2 0 740 1 1 1 0
17 299 0 0 1 601 1 2 399 640 2 1 0 0
Table 2: The $M=17$-run CA design and test outcomes for our motivating TCAS case study. Here, an outcome of 0 indicates a passed test case, with 1 indicating a failed test case.
2.2 State-of-the-Art and Its Limitations
Existing fault localization methods, as described earlier, are largely deterministic in nature. This includes the work of Niu et al., (2013), who used a tuple relationship tree to capture relationships among all factor combinations based on testing results. For a given test case, this tree lists all considered combinations (tuples) along the branches of the tree, which can then be used to classify which class (faulty or healthy) each tuple belongs to. Nie and Leung, 2011a introduced the idea of a minimal failure-causing schema, defined as the smallest-order factor combination such that all test cases containing them trigger a failure. This schema is then applied for guiding fault localization and subsequent testing. Ghandehari, (2016) and Ghandehari et al., (2018) developed a two-stage combinatorial-testing-based fault localization approach. The key idea is to identify potential failure-inducing factor combinations from test results by eliminating combinations that appear in passed test cases, then rank such combinations based on two proposed “combination suspiciousness” and “environment suspiciousness” metrics. Lekivetz and Morgan, (2018) proposed a deterministic ranking procedure that incorporates a domain-knowledge-guided weighting scheme. A key limitation of such methods is that they are not probabilistic in nature, and thus do not provide the desired probabilistic measure of confidence for how likely a particular combination is a root cause. The above methods also unfortunately do not have publicly available code; we instead make use of the JMP Covering Array Analysis (JMP Statistical Discovery LLC,, 2023) as the “state-of-the-art” approach, which has integrated such methods either in full or in part.
We first provide a brief overview of the fault localization analysis from the JMP Covering Array module. Similar to Ghandehari et al., (2018), this analysis first removes all combinations that have been cleared in passed cases; we call these “tested-and-passed” combinations later. The remaining combinations are then ranked in terms of its failure count, i.e., the number of failed test cases for which this combination is present. For combinations with the same failure count, those with the smallest-order combination are treated as more suspicious, following the principle of combination hierarchy in Lekivetz and Morgan, (2018). Here, the ranking of potential root cause combinations via its failure count is quite intuitive, as combinations that show up in more failed test cases should naturally be treated as more suspicious.
For our TCAS case study, Figure 2 shows the top ten combinations (all tied in terms of suspiciousness) returned by the JMP Covering Array Analysis. This analysis highlights several limitations with the state-of-the-art. First, the tied nature of these top ten combinations makes its subsequent investigation very costly. For such complex systems, the debugging of each combination can require thousands of dollars (Bush,, 1990); the budget required for debugging ten combinations can thus be prohibitively expensive. Second, this analysis is deterministic and does not shed light on the probability that a suspicious combination is a root cause. Such a probabilistic quantification of risk is needed for trustworthy fault diagnosis (Zhou et al.,, 2023). A further benefit of a probabilistic analysis is its potential for disentangling the tied combinations noted above. This potential for disentangling effects was noted in Chang, (2019), who observed that a Bayesian framework can help de-alias effects in fractional factorial designs; see also related work by Mak and Wu, (2019). We show later that a carefully-specified Bayesian model can similarly help disentangle the many tied combinations from Figure 2, facilitating timely and cost-efficient fault diagnosis.
Figure 2: Top suspicious combinations from the JMP Covering Array analysis for our motivating TCAS case study, with its corresponding failure counts from test runs.
3 The BayesFLo Model
We propose a new Bayesian Fault Localization (BayesFLo) framework, which provides a principled probabilistic approach for fault localization via conditioning on test set outcomes. We first present the employed Bayesian model, then show how it permits the integration of domain and structural knowledge within its prior specification. With such knowledge integrated, BayesFLo can then facilitate timely and confident fault localization with limited costly test runs.
Here, it is worth mentioning that, while there may be different “types” of failures (e.g., incorrect or missing outputs), the system has typically undergone a reasonable level of preliminary testing (e.g., unit testing, which eliminates root causes in single units) at time of combinatorial testing. Because of this, there remains only a very limited number of faults for diagnosis (Lekivetz and Morgan,, 2021). As such, we presume in the following only a single failure type at time of BayesFLo analysis. We will briefly discuss an approach for the rare scenario of multiple failure types later in the Conclusion.
3.1 Prior Specification
We first introduce some notation. Consider a software system (or more broadly, a complex engineering system) with $I\geq 1$ input factors, where a factor $i$ can take on $J_{i}\geq 2$ different levels. A $K$-input combination (with $K\leq I$) is denoted as $(\mathbf{i},\mathbf{j})_{K}$, where $\mathbf{i}=(i_{1},\cdots,i_{K}),i_{1}<\cdots<i_{K}$ is an ordered $K$-vector containing all inputs for this combination, and $\mathbf{j}=(j_{1},\cdots,j_{K}),j_{k}\in\{1,\cdots,J_{i_{k}}\}$ is a $K$-vector indicating the levels of each corresponding input. For example, the 2-input combination of the first factor at level 1 and the second factor at level 2 can be denoted as $(\mathbf{i},\mathbf{j})_{2}$, where $\mathbf{i}=(1,2)$ and $\mathbf{j}=(1,2)$. In the case of $K=1$, i.e., a single input $i$ at level $j$, this may be simplified to $(i,j)$.
Now let $\mathcal{C}_{K}$ denote the set of $K$-input combinations $(\mathbf{i},\mathbf{j})_{K}$ as described above, and let $\mathcal{C}=\cup_{K=1}^{I}\mathcal{C}_{K}$ denote the set of combinations over all orders $K=1,\cdots,I$. Further let $Z_{(\mathbf{i},\mathbf{j})_{K}}\in\{0,1\}$ be an indicator variable for whether the combination $(\mathbf{i},\mathbf{j})_{K}$ is truly a root cause. As this is unknown prior to running test cases, we model each $Z_{(\mathbf{i},\mathbf{j})_{K}}$ a priori as an independent Bernoulli random variable of the form:
$Z_{(\mathbf{i},\mathbf{j})_{K}}\overset{indep.}{\sim}\text{Bern}\{p_{(\mathbf{i},\mathbf{j})_{K}}\},\quad(\mathbf{i},\mathbf{j})_{K}\in\mathcal{C}_{K},\quad K=1,\cdots,I,$ (1)
where $p_{(\mathbf{i},\mathbf{j})_{K}}$ is the prior probability that this combination is a root cause. Here, the view that $Z_{(\mathbf{i},\mathbf{j})_{K}}$ is random makes our approach Bayesian; this contrasts with existing fault localization approaches, which presume $Z_{(\mathbf{i},\mathbf{j})_{K}}$ to be fixed but unknown. For $K=1$, this notation simplifies to $Z_{(i,j)}$ and $p_{(i,j)}$. Whenever appropriate, we denote $\mathbf{Z}=(Z_{(\mathbf{i},\mathbf{j})_{K}})_{(\mathbf{i},\mathbf{j})_{K}\in\mathcal{C}}$ and $\mathbf{p}=(p_{(\mathbf{i},\mathbf{j})_{K}})_{(\mathbf{i},\mathbf{j})_{K}\in\mathcal{C}}$ for brevity.
It is worth noting the sheer number of input combinations in $\mathcal{C}$ that needs to be considered as potential root causes. Assuming each factor has an equal number of levels $J=J_{1}=\cdots=J_{I}$, one can show that $\mathcal{C}_{K}$ contains $\binom{I}{K}J^{K}$ distinct combinations of order $K$, thus the total number of consider input combinations is $|\mathcal{C}|=$ $\sum_{K=1}^{I}\binom{I}{K}J^{K}$. Even with a moderate number of inputs, say $I=10$, with each having $J=2$ levels, this amounts to $|\mathcal{C}|=59,048$ combinations. As we shall see later in Section 4, the size of $\mathcal{C}$ forms the key bottleneck for Bayesian inference, as the computation of posterior probabilities can require $\mathcal{O}(2^{|C|})$ work; this can thus be infeasible even for small software systems.
Next, we adopt the following product form on the root cause probability for $(\mathbf{i},\mathbf{j})_{K}$:
$p_{(\mathbf{i},\mathbf{j})_{K}}=\prod_{k=1}^{K}p_{(i_{k},j_{k})},\quad(\mathbf{i},\mathbf{j})_{K}\in\mathcal{C}_{K},\quad K=2,\cdots,I.$ (2)
In words, the combination root cause probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ is modeled as the product of the root cause probabilities for its component inputs. The key advantage of this form is that it captures structural knowledge on root causes, by embedding the principles of combination hierarchy and heredity (Lekivetz and Morgan,, 2021). These principles capture the structured nature of typical software root causes, and can be seen as extensions of the well-known principles of effect hierarchy and heredity (Wu and Hamada,, 2009), which are widely used for analysis of factorial experiments. The first principle, combination hierarchy, asserts that combinations involving fewer inputs are more likely to be failure-inducing than those involving more inputs. Empirical evidence suggests this principle holds for software systems across various domains (Kuhn et al.,, 2004). To see how our prior (2) captures combination hierarchy, note that by its product form construction, the combination probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ is always less than the probability of any component input $p_{(i_{k},j_{k})}$. Thus, this prior assigns increasingly smaller root cause probabilities on combinations with a higher combination order $K$, thus capturing the desired hierarchy structure. The second principle, combination heredity, asserts that a combination is more likely to be failure-inducing when some of its component inputs are more likely to be failure-inducing. From our product-form prior (2), note that the combination root cause probability $p_{(\mathbf{i},\mathbf{j})_{K}}$ cannot be large unless some of its component root cause probabilities in $\{p_{(i_{k},j_{k})}\}_{k=1}^{K}$ are also large. This thus captures the desired combination heredity effect. Similar product-form weights have been used for modeling hierarchy and heredity in the context of predictive modeling (Tang et al.,, 2023) and data reduction (Mak and Joseph,, 2017).
With the product-form prior (2), we require only the specification of the single-input root cause probabilities $\{p_{(i,j)}\}_{i,j}$, which can be used to integrate further domain knowledge from test engineers. For most software systems at the testing stage, it may be reasonable to specify a small (i.e., near-zero) value for $p_{(i,j)}$, as this reflects the prior belief that failure-inducing root causes should occur sporadically. Oftentimes, however, an engineer has additional domain knowledge that permits a more informed prior specification, e.g., from historical data or expert judgment. For example, the engineer may know that certain factors have been recently added to the system, and thus may be more suspicious of such factors. This heightened suspicion can be captured via a larger specification of its $p_{(i,j)}$ compared to other factors. We shall see how such domain knowledge can accelerate fault localization in later case studies.
From a Bayesian perspective, the product-form prior (2) provides a way for propagating elicited domain knowledge over the many root cause probabilities in $\mathbf{p}$. For example, suppose an engineer has heightened suspicions on factor $i$, and accordingly specifies a higher value for the single-factor root cause probabilities $\{p_{(i,j)}\}_{j}$. By (2), this induces larger prior root cause probabilities $p_{(\mathbf{i},\mathbf{j})_{K}}$ for any combination $(\mathbf{i},\mathbf{j})_{K}$ involving factor $i$, thus “pooling” this information over such combinations. This “information pooling”, guided by the embedded principles of combination hierarchy and heredity, can facilitate the disentangling of the large number of potential root causes from limited test runs. Recent work on related notions of information pooling have shown promise in high-dimensional inference problems, e.g., matrix completion (Yuchi et al.,, 2023) and multi-armed bandits (Mak et al.,, 2022); we show that this is also similarly important for timely fault localization. Of course, as with any Bayesian modeling framework, the domain and/or structural knowledge captured within the BayesFLo prior is not a strict assumption; it can be overturned with sufficient evidence from data.
3.2 Posterior Root Cause Probabilities
In what follows, we suppress the notation $(\mathbf{i},\mathbf{j})_{K}$ to $(\mathbf{i},\mathbf{j})$ for brevity. Using the above prior specification, we now need to condition on the test outcome data. Suppose we run the software system at $M$ different test cases, where the $m$-th test case is performed at input levels $\mathbf{t}_{m}=(t_{m,1},\cdots,t_{m,I})$, $t_{m,i}\in\{1,\cdots,J_{i}\}$. Then the test data can be denoted as $\mathcal{D}=\{(\mathbf{t}_{m},y_{m})\}_{m=1}^{M}$, where $y_{m}\in\{0,1\}$ is a binary variable with 1 indicating a failure and 0 if not. To make things concrete, consider the following example. Suppose the system has $I=3$ input factors, each with two levels. Further assume there is only one true root cause $((1,2),(1,2))$, i.e., the combination of the first input at level 1 and the second at level 2, which results in failure. Suppose we then run the first test case at input setting $\mathbf{t}_{1}=(1,2,1)$, i.e., with the three factors at levels 1, 2 and 1, respectively. Then, since the root cause is present in $\mathbf{t}_{1}$, this would result in a failure, namely $y_{1}=1$. However, if we run the second test case at a different setting $\mathbf{t}_{2}=(2,2,2)$, then this test case would result in no failure, i.e., $y_{2}=0$, as the root cause is not present in $\mathbf{t}_{2}$. Here, we presume that observed outcomes are deterministic, in that the same outcome $y_{m}$ is always observed whenever the software system is run with inputs $\mathbf{t}_{m}$.
With this framework, the problem of fault localization then reduces to the evaluation of the posterior root cause probabilities for all considered combinations in $\mathcal{C}$, namely:
$\small\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D}),\quad\text{for all}\;(\mathbf{i},\mathbf{j})\in\mathcal{C}.$ (3)
Such a computation, however, can easily become computationally intractable. The key bottleneck lies in the complex conditioning structure from data $\mathcal{D}$ over the high-dimensional set of combinations $\mathcal{C}$; as we see later, this can then induce an $\mathcal{O}(2^{|\mathcal{C}|})$ complexity for a brute-force computation of posterior probabilities. Recall that, with the moderate setting of $I=10$ and $J=2$, $|\mathcal{C}|$ consists of nearly 60,000 combinations. Thus, without careful modifications to exploit problem structure, posterior computation can be intractable even for small systems.
We adopt next the following categorization of input combinations in $\mathcal{C}$ for efficient computation of root cause probabilities:
  1. (a)
    Tested-and-Passed (TP): TP combinations for a passed test case $\mathbf{t}_{m}$ are combinations in $\mathcal{C}$ that have been tested in $\mathbf{t}_{m}$. Continuing from the earlier example, suppose we run the test case $\mathbf{t}_{m}=(2,2,2)$ with no failure, i.e., with $y_{m}=0$. Then it follows that the combination $((1,2),(2,2))$, i.e., with the first factor A at level 2 and the second factor B at level 2, is a TP combination. (In what follows, we may denote such a combination as $\text{A}_{2}\text{B}_{2}$ for notational simplicity; this should be clear from context.) For this single passed case, the set of TP combinations is $\mathcal{C}_{{\rm TP},m}=\{\text{A}_{2},\text{B}_{2},\text{C}_{2},\text{A}_{2}\text{B}_{2},\text{A}_{2}\text{C}_{2},\text{B}_{2}\text{C}_{2},\text{A}_{2}\text{B}_{2}\text{C}_{2}\}$.
  2. (b)
    Tested-and-Failed (TF): TF combinations for a failed test case $\mathbf{t}_{m}$ are combinations in $\mathcal{C}$ that have been tested in $\mathbf{t}_{m}$. For example, suppose we run the test case $\mathbf{t}=(1,2,1)$ and observe a failure, i.e., with $y_{m}=1$. Then, from this single failed case, the set of TF combinations becomes $\mathcal{C}_{{\rm TF},m}=\{\text{A}_{1},\text{B}_{2},\text{C}_{1},\text{A}_{1}\text{B}_{2},\text{A}_{1}\text{C}_{1},\text{B}_{2}\text{C}_{1},\text{A}_{1}\text{B}_{2}\text{C}_{1}\}$.
  3. (c)
    Untested (UT): UT combinations are combinations in $\mathcal{C}$ that have not been tested in any test case. For example, suppose we run the test case $\mathbf{t}=(1,2,1)$. Then one UT combination is $\text{A}_{2}\text{B}_{1}$, as such a combination was not tested in $\mathbf{t}$.
This partition of $\mathcal{C}$ naturally extends for multiple test runs in $\mathcal{D}$. Here, the TP combinations $\mathcal{C}_{\rm TP}$ from $\mathcal{D}$ are the TP combinations over all passed test cases. The TF combinations $\mathcal{C}_{\rm TF}$ from $\mathcal{D}$ are the TF combinations over all failed test cases, with the combinations from $\mathcal{C}_{\rm TP}$ removed. $\mathcal{C}_{\rm UT}$ then consists of all remaining combinations in $\mathcal{C}$. In other words:
$\mathcal{C}_{\rm TP}=\cup_{m:y_{m}=0}\;\mathcal{C}_{{\rm TP},m},\quad\mathcal{C}_{\rm TF}=(\cup_{m:y_{m}=1}\;\mathcal{C}_{{\rm TF},m})\setminus\mathcal{C}_{\rm TP},\quad\mathcal{C}_{\rm UT}=\mathcal{C}\setminus(\mathcal{C}_{\rm TP}\cup\mathcal{C}_{\rm TF}).$ (4)
Figure 3 (left) visualizes this partition of $\mathcal{C}$ from observed test runs for a simple example.
Figure 3: [Left] Visualizing the use of passed and failed test cases for partitioning the set of considered combinations $\mathcal{C}$ into $\mathcal{C}_{\rm TP}$, $\mathcal{C}_{\rm TF}$ and $\mathcal{C}_{\rm UT}$. [Right] Workflow for the proposed BayesFLo fault localization approach.
With this partition of $\mathcal{C}$, we now investigate efficient algorithms for computing the posterior root cause probabilities (3). For TP combinations, it is clear that such a combination cannot be a root cause, as it was cleared by a passed test case. In other words:
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})=0,\quad(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TP}.$ (5)
This is akin to Ghandehari et al., (2018), which removes TP combinations from consideration for root causes. Furthermore, for UT combinations, we have:
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})=\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1),\quad(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm UT},$ (6)
since the observed test set $\mathcal{D}$ does not provide any information on an untested combination $(\mathbf{i},\mathbf{j})$. As such, its root cause probability given $\mathcal{D}$ simply reduces to its prior probability given in (2). The challenge thus lies in computing posterior probabilities on the remaining class of TF combinations. We present next an approach for computing such probabilities, leveraging tools from integer programming and graph representations. Figure 3 (right) summarizes the proposed algorithmic workflow; we elaborate on this in the following section.
4 Computation of Root Cause Probabilities
Consider the case of TF combinations, where we wish to compute the posterior root cause probability (3) for a given TF combination $(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TF}$. One solution might be the following “brute-force” approach:
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})=\frac{\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1,\mathcal{D})}{\mathbb{P}(\mathcal{D})}=\frac{\sum_{\mathbf{z}\in\{0,1\}^{|\mathcal{C}|},Z_{(\mathbf{i},\mathbf{j})}=1}\mathbb{P}(\mathbf{Z}=\mathbf{z})\mathbb{P}(\mathcal{D}|\mathbf{Z}=\mathbf{z})}{\sum_{\mathbf{z}\in\{0,1\}^{|\mathcal{C}|}}\mathbb{P}(\mathbf{Z}=\mathbf{z})\mathbb{P}(\mathcal{D}|\mathbf{Z}=\mathbf{z})}.$ (7)
where $\mathbb{P}(\mathbf{Z}=\mathbf{z})$ follows from Equation (2), and $\mathbb{P}(\mathcal{D}|\mathbf{Z}=\mathbf{z})$ can be deduced from the reasoning in Section 3.2. The limitation of such an approach is clear. For each $(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TF}$, we need to compute the sum of $2^{|\mathcal{C}|-1}$ terms in the numerator and the sum of $2^{|\mathcal{C}|}$ terms in the denominator. Hence, even for small systems with $|\mathcal{C}|$ small, this brute-force approach can be infeasible. This sheer dimensionality of potential root cause scenarios is the key bottleneck for tractable computation of probabilities for Bayesian fault localization.
To address this, we employ an alternate formulation, which permits considerable speed-ups for computing probabilities. We first outline this reformulation, then show how this facilitates efficient computation via a connection to the related problem of minimal set covering.
4.1 An Alternate Formulation
The following proposition provides a useful reformulation of the desired posterior root cause probability for a TF combination $(\mathbf{i},\mathbf{j})$:
Proposition 1.
Let $(\mathbf{i},\mathbf{j})\in\mathcal{C}_{\rm TF}$, and let:
$\mathcal{M}_{(\mathbf{i},\mathbf{j})}=\{m=1,\cdots,M:y_{m}=1,(\mathbf{i},\mathbf{j})\in\mathcal{C}_{{\rm TF},m}\}$ (8)
be the index set of failed test cases for which $(\mathbf{i},\mathbf{j})$ is a potential root cause. Define the event:
$E_{(\mathbf{i},\mathbf{j})}=\{\textup{for each $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m}\setminus\mathcal{C}_{{\rm TP}}$ such that $Z_{c}=1$}\}.$ (9)
In words, this is the event that all failures in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ can be explained by the selected root causes $\{c\in\mathcal{C}_{\rm TF}:Z_{c}=1\}$. The desired posterior root cause probability then reduces to:
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})=\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|E_{(\mathbf{i},\mathbf{j})})=\frac{p_{(\mathbf{i},\mathbf{j})}}{\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})}.$ (10)
The proof of this proposition can be found in Appendix A. There are two key advantages of the alternate form (10) over the brute-force approach (7). First, its numerator can be directly computed via Equation (2) with little work. Second, its denominator $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ can be effectively computed via a novel connection to a related minimal set covering problem for bipartite graphs (Asratian et al.,, 1998), which we show below.
To compute $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$, we first inspect condition (9) for $E_{(\mathbf{i},\mathbf{j})}$, which requires, for each failed test case in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, a corresponding TF combination that induces this failure. Figure 4 visualizes this condition in the form of a bipartite graph, where the left nodes are the TF combinations in $\mathcal{C}_{\rm TF}$, and the right nodes are failed test cases in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$. Here, an edge is drawn from a combination $c$ (on left) to a test case index $m$ (on right) if $c\in\mathcal{C}_{{\rm TF},m}$, i.e., if combination $c$ is contained in the failed test inputs $\mathbf{t}_{m}$. Viewed this way, condition (9) is equivalent to finding a selection of potential root causes in $\{Z_{c}\}_{c\in\mathcal{C}_{\rm TF}}$, such that every failed test case on the right is connected to (or “covered” by) a selected combination on the left via an edge. Figure 4 visualizes two possible “covers”. Such a cover of right-hand nodes can be interpreted as a selection of potential root causes (left-hand nodes) that explains the failed test cases. Thus, to compute the probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$, we need to sum over the prior probabilities for all possible selections of potential root causes that cover the failed test cases in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$.
4.2 Enumerating Minimal Covers
Using this insight, we now establish a useful link between the desired probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ and the related problem of minimal set covering. Formally, we define a cover of the failed test indices $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ as a subset $\tilde{\mathcal{C}}$ of the potential root causes $\mathcal{C}_{\rm TF}$, such that for every $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists an edge connecting some node in $\tilde{\mathcal{C}}$ to $m$. A minimal cover of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ is then a cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ which, if any element is removed from $\tilde{\mathcal{C}}$, ceases to be a cover. Figure 4 visualizes this notion of a minimal cover.
Figure 4: Visualizing the bipartite graph representation and two minimal covers for failures involving the combination $(\mathbf{i},\mathbf{j})$.
With this definition, the following proposition reveals a useful connection:
Proposition 2.
The desired probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ can be simplified as:
$\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})=\mathbb{P}(\{Z_{c}=1\textup{ for all }c\in\tilde{\mathcal{C}}\}\textup{, for at least one minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$}).$ (11)
Its proof can be found in Appendix B. This shows that $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ is equivalent to the probability that, for at least one minimal cover $\tilde{\mathcal{C}}$, all combinations in $\tilde{\mathcal{C}}$ are indeed root causes.
To compute (11), a natural approach is to first enumerate all minimal covers of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$. Fortunately, the set cover problem for bipartite graphs has been well-studied in the literature, and efficient polynomial-time algorithms have been developed for finding minimal covers (Skiena,, 1998; Hopcroft and Karp,, 1973). Leveraging such developments can thus greatly speed up the brute-force approach for posterior probability computation (see Equation (7)), which is doubly-exponential in complexity and thus infeasible for even small software systems. With recent developments in integer programming algorithms (Wolsey,, 2020), a popular strategy for finding minimal set covers is to formulate and solve this problem as an integer linear program (ILP; Schrijver,, 1998). We adopt such a strategy below.
Let $\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}$ be the set of potential root causes in $\mathcal{C}_{\rm TF}$ involving $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$; this is typically much smaller than $\mathcal{C}_{\rm TF}$, which reduces the size of the optimization program below. We propose the following feasibility program to find one minimal cover for $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$:
$\begin{split}\operatorname*{arg\,max}\;1&\text{\quad s.t. \quad}\\ &z_{c}\in\{0,1\}\;\text{for all }c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},\quad l_{g,m}\in\{0,1\}\;\text{for all }g\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})},\\ \text{[C1] }&\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}}z_{c}\cdot\mathbb{I}(m\in\mathcal{M}_{c})\geq 1\;\text{for all }m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})},\\ \text{[C2] }&\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},c\neq g}z_{c}\cdot\mathbb{I}(m\in\mathcal{M}_{c})\leq|\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}|(1-l_{g,m})\;\text{for all }g\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})},\\ \text{[C3] }&\sum_{m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}}l_{g,m}\geq 1\;\text{for all }g\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}.\end{split}$
The decision variables are the binary variables $\{z_{c}\}$ and $\{l_{g,m}\}$, with $z_{c}=1$ indicating combination $c$ is included in the cover and $z_{c}=0$ otherwise. Here, the dummy objective of 1 is used, as we only wish to find a feasible solution satisfying constraints [C1]-[C3]. The first constraint [C1] requires the selected combinations $\{c:z_{c}=1\}$ to cover all failed test cases in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$. The next constraints [C2] and [C3] ensure the selected cover is indeed a minimal cover. To see why, note that via constraint [C2], the auxiliary indicator variable $l_{g,m}\in\{0,1\}$ equals 1 if by removing $g$ from the considered cover, we fail to cover failure case $m$. For the considered cover to be minimal, we thus need, for each $g$ in the cover, at least one $l_{g,m}=1$ for some failure case $m$; this is ensured by constraint [C3].
One appealing property of the integer feasible program (12) is that the objective is (trivially) linear and all constraints are linear in the binary decision variables. Such an integer linear program thus admits nice structure for efficient large-scale optimization, particularly via recent developments in cutting plane and branch-and-bound algorithms (Balas et al.,, 1993; Stidsen et al.,, 2014). In our later implementation, we made use of the GurobiPy package in Python (Gurobi Optimization, LLC,, 2023), which implements state-of-the-art optimization solvers for large-scale integer programming. Gurobi is widely used for solving large-scale scheduling problems in the industry, including for the National Football League (North,, 2020) and Air France (Richard,, 2020). Here, with the ILP formulation (12), Gurobi can solve for a feasible minimal set cover in minutes for our later case studies. This formulation thus provides an efficient strategy for computing the desired probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$.
Of course, after finding a single minimal cover via (12), we still have to find subsequent distinct minimal covers to compute (11). This can easily be performed by iteratively solving the ILP (12) with an additional constraint that ensures subsequent covers are distinct from found covers. More concretely, let $\{\tilde{z}_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$ be a minimal cover found by (12). Then a subsequent cover can be found by solving the ILP (12) with the additional constraint:
$d_{c}=z_{c}\oplus\tilde{z}_{c},\quad\sum_{c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})}}d_{c}\geq 1,\quad c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},$ [C4]
where $\oplus$ is the XOR operator. This new constraint [C4] ensures the next cover is distinct from the previous found cover. To see why, note that $d_{c}$ equals 1 only if the binary variables $z_{c}$ and $\tilde{z}_{c}$ are different; the inequality constraint in [C4] thus ensures all considered covers are different from the previous cover $\{\tilde{z}_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$. The resulting problem is still an integer linear program, as XOR can naturally be expressed as linear constraints (Magee and Glover,, 1996). More specifically, the XOR condition in [C4] can be equivalently expressed as:
$\begin{split}d_{c}\geq z_{c}-\tilde{z}_{c},\quad d_{c}\geq\tilde{z}_{c}-z_{c},\quad d_{c}\leq z_{c}+\tilde{z}_{c},\quad d_{c}\leq 2-z_{c}-\tilde{z}_{c},\quad c\in\mathcal{C}_{\text{TF},(\mathbf{i},\mathbf{j})},\end{split}$ (13)
which is clearly linear in the binary decision variables $\{z_{c}\}_{c\in\mathcal{C}_{\rm TF,(\mathbf{i},\mathbf{j})}}$.
With this, if a feasible solution is found for the ILP (12) with [C4], the optimization solver will return a distinct minimal cover, which we add to the collection. If not, the solver will instead return a “dual certificate” (Güzelsoy et al.,, 2010) that guarantees the ILP has no feasible solutions; such a certificate is possible via the linear nature of the integer program. One then iteratively solves the ILP (12) with constraint [C4] (modified to exclude two or more found covers) until the solver returns a dual certificate, in which case no feasible solutions are possible and thus all minimal covers have been enumerated.
4.3 Computing Root Cause Probabilities
After enumerating all minimal covers for $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, we can then compute $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ via Proposition 11. Let $\mathcal{V}=\{\tilde{\mathcal{C}}_{1},\cdots,\tilde{\mathcal{C}}_{|\mathcal{V}|}\}$ be the collection of all minimal covers of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ found by the above procedure. By the principle of inclusion-exclusion, it follows from (11) that:
$\begin{split}\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})&=\mathbb{P}(\{Z_{c}=1\textup{ for all }c\in\tilde{\mathcal{C}}\}\textup{, for at least one $\tilde{\mathcal{C}}\in\mathcal{V}$})\\ &=\sum_{\text{cover }\tilde{\mathcal{C}}\in\mathcal{V}}\;\prod_{c\in\tilde{\mathcal{C}}}p_{c}-\sum_{\text{covers }\tilde{\mathcal{C}},\tilde{\mathcal{C}}^{\prime}\in\mathcal{V}}\;\prod_{c\in\tilde{\mathcal{C}}\cup\tilde{\mathcal{C}}^{\prime}}p_{c}+\cdots+(-1)^{|\mathcal{V}|}\prod_{c\in{\tilde{\mathcal{C}}_{1}\cup\cdots\cup\tilde{\mathcal{C}}_{|\mathcal{V}|}}}p_{c},\end{split}$
where $p_{c}=\mathbb{P}(Z_{c}=1)$ is again the prior root cause probability of combination $c$. We can then plug the computed $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$ into Equation (10) to finally compute the desired root cause probability $\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})$ for a TF combination $(\mathbf{i},\mathbf{j})$.
For software systems with a small number of inputs, the set of minimal covers $\mathcal{V}$ may not be large, in which case the computation in (14) would not be intensive. For larger systems with $|\mathcal{V}|$ large, one can employ the following second-order truncation as an approximation:
$\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})\approx\sum_{\text{cover }\tilde{\mathcal{C}}\in\mathcal{V}}\;\prod_{c\in\tilde{\mathcal{C}}}p_{c}-\sum_{\text{covers }\tilde{\mathcal{C}},\tilde{\mathcal{C}}^{\prime}\in\mathcal{V}}\;\prod_{c\in\tilde{\mathcal{C}}\cup\tilde{\mathcal{C}}^{\prime}}p_{c},$ (15)
which bypasses the need for computing higher-order terms involving more than two covers. Note that, by the inclusion-exclusion principle, the right-hand side of (15) underestimates the probability $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})$. This is by design: from (10), this then results in a slight overestimation of the posterior root cause probability $\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})$. From a risk perspective, this is more preferable than an approximation procedure that underestimates such probabilities.
4.4 Algorithm Summary
For completeness, we provide in Algorithm 1 a summary of the full BayesFLo procedure. Suppose a test set is performed, yielding outcome data $\mathcal{D}=\{(\mathbf{t}_{m},y_{m})\}_{m=1}^{M}$. Here, the test cases $\{\mathbf{t}_{m}\}_{m=1}^{M}$ should ideally be collected from a covering array to ensure good coverage of combinations, but this is not necessary for BayesFLo. With test data collected and priors elicited on the single-factor root cause probabilities $\{p_{(i,j)}\}_{i,j}$, we then partition the set of considered combinations $\mathcal{C}$ into TP, TF and UT combinations using Equation (4). Here, if the test engineer is confident that a root cause should not exceed a certain order, then posterior probabilities need only to be computed for combinations up to such an order. This can be justified as a stronger form of combination hierarchy, and can further reduce computation for evaluating posterior probabilities.
Next, we compute posterior root cause probabilities within each category. For TP combinations, this is trivially zero as such combinations were cleared in passed cases. For UT combinations, this can be set as the prior probabilities from (2), as no information can be gleaned on such combinations from the test data. For TF combinations, its posterior probabilities can be computed via the minimal set cover approach in Section 4. Finally, with posterior probabilities computed, we can then rank the potential root causes in terms of their probabilities, which can be used for guiding software diagnosis; more on this later. Figure 3 (right) visualizes the full workflow for the BayesFLo procedure.
We conclude this section with a discussion of computation time for BayesFLo. When the test data consists of only a single failed test run, the BayesFLo procedure from Algorithm 1 runs very quickly in our experience. For our later Case Study 2, where only one failed test run is observed (see Table 5), all posterior BayesFLo probabilities can be computed in several seconds, which is comparable in time to the state-of-the-art JMP Covering Analysis. When the test data consists of multiple failed runs, BayesFLo can require considerably more computation time, particularly for minimal cover enumeration. For our later Case Study 1, where multiple failed runs are observed (see Table 2), posterior BayesFLo probabilities require several hours of computation, whereas the JMP Covering Analysis takes only seconds.
This increased computation time for BayesFLo in the case of multiple failed test runs is, however, not a major bottleneck for three reasons. First, the key advantage of BayesFLo over the state-of-the-art lies in its improved probabilistic analysis, which permits confident fault localization with a small number of costly test runs. Thus, while our analysis may require more computation, this is in many cases preferable to committing tens of thousands of dollars on investigating many tied suspicious combinations, as shown in later case studies. Second, as alluded to earlier, we presume that at time of performing BayesFLo, the system has been sufficiently tested such that few failed runs are expected. Finally, with the increasing availability of multi-core processors, the computation for BayesFLo can be considerably accelerated by exploiting such parallelism within the GurobiPy package; see Glockner, (2015) for details.
Algorithm 1 BayesFLo: Bayesian Fault Localization
{justify}
Input: Test data $\mathcal{D}=\{(\mathbf{t}_{m},y_{m})\}_{m=1}^{M}$, consisting of each test case and its corresponding test outcome.
Output: Potential root causes $(\mathbf{i},\mathbf{j})\in\mathcal{C}$ with posterior root cause probabilities $\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})$.
1:Elicit root cause probabilities $\{p_{(i,j)}\}_{i,j}$ from domain knowledge.
2:Partition the set of considered combinations $\mathcal{C}$ into TP, TF and UT combinations using Equation (4).
3:For TP combinations, set its posterior root cause probability to 0.
4:For UT combinations, set its posterior root cause probability as the prior probability (2).
5:For each TF combination $(\mathbf{i},\mathbf{j})$, enumerate minimal covers for the failed cases in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, then compute its posterior root cause probability using Equation (15).
6:Rank potential root causes (TF and UT combinations) using its corresponding posterior probabilities.
5 Case Studies
We now explore the effectiveness of BayesFLo for tackling our motivating fault localization case study on the Traffic Alert and Collision Avoidance System (Section 2), as well as another case study on the JMP Easy DOE platform (Jones and Sall,, 2011). The first case study illustrates the importance of a flexible Bayesian modeling framework that integrates structural knowledge for disentangling the many tied combinations from existing analyses. The second explores the importance of integrating domain knowledge to disentangle tied combinations for confident fault localization.
5.1 Case Study 1: Traffic Alert and Collision Avoidance System
Consider first the motivating case study in Section 2, on the fault localization of the Traffic alert and Collision Avoidance System (TCAS). As mentioned earlier, TCAS targets the avoidance of potential mid-air collisions between aircraft, and its reliable performance is essential for mitigating catastrophic disasters (see, e.g., Davies and Yousif,, 2025). In what follows, we consider a specific program – the Siemens test suite in Griesmayer et al., (2007) – for fault localization. Its $I=12$ input parameters with corresponding levels are summarized in Table 1. We consider the “TCAS v2” set-up from Griesmayer et al., (2007), which contains a fully-documented fault that results in an incorrect resolution advisory; if unaddressed, such a fault may lead to potential aircraft collisions. Here, it is known (Griesmayer et al.,, 2007) that such a failure arises from a single root cause, namely, the three-factor combination of factors 12, 8 and 9:
[…]
With this set-up, we then construct a strength-2 covering array with the input factors and levels from Table 1, using a run size of $M=17$. Table 2 shows the corresponding test outcomes for each test run. Figure 2 outlines the fault localization analysis from the JMP Covering Array module, which we treat as the state-of-the-art approach that integrates several existing methods. Recall that such an analysis has several limitations. First, its top ten combinations are all tied in terms of suspiciousness (i.e., its failure count); the budget needed to investigate all ten combinations can thus be prohibitively expensive, as the debugging of each combination can require thousands of dollars (Bush,, 1990). Second, this analysis is deterministic and offers no insight on the probability that a suspicious combination is indeed a root cause given observed test outcomes. Such probabilistic analysis facilitates confident fault localization and permits the disentangling of tied combinations for timely diagnosis.
To implement BayesFLo, we first need a careful specification of its prior distribution in (2). Here, we have little prior domain knowledge besides the belief that root causes are rare. To reflect this, we adopt the weakly-informative prior of $p_{(i,j)}=1/30$, where the denominator $\sum_{i=1}^{I}J_{i}=30$ is the total number of levels for all $I=12$ factors. In what follows, we consider root cause combinations with at most three factors; this not only reduces computation, but is also guided by prior observations that root causes are typically of lower order for such systems (Kuhn et al.,, 2004). Posterior root cause probabilities are then computed via Algorithm 1.
Top Suspicious Combinations (BayesFLo)
Factors Levels Posterior Probability
12, 8, 9 1, 399, 640 0.55
8, 7 399, 1 0.20
9, 7 640, 1 0.20
7, 10 1, 0 0.20
9, 1 640, 300 0.10
9, 11 640, 2 0.10
7, 10 0, 1 0.10
7, 2 0, 0 0.10
7, 6 0, 2 0.10
1, 10 300, 1 0.10
10, 11 1, 2 0.10
Table 3: Top suspicious combinations from the BayesFLo analysis in the TCAS case study, ranked by decreasing posterior root cause probabilities. Remaining combinations all have posterior probabilities less than $7.0\times 10^{-3}$. Here, the true root cause is bolded.
Table 3 shows the top suspicious combinations from the BayesFLo analysis, ranked by its posterior root cause probabilities. There are two key observations to note. First, compared to the JMP analysis (Figure 2), BayesFLo returns a more informative list that disentangles tied combinations: it identifies the three-factor combination (bolded in Table 3) as the most suspicious with a posterior probability of 0.55, followed by three two-factor combinations with posterior probabilities of 0.20. This top three-factor combination from BayesFLo is precisely the true root cause (16); test engineers can thus quickly identify the true root cause from our analysis, which greatly reduces costs. This highlights the importance of a flexible probabilistic modeling framework that embeds structural principles for disentangling many tied combinations. Second, BayesFLo provides a probabilistic quantification of combination suspiciousness via its posterior root cause probabilities, including for higher-order combinations that would not have been considered by existing methods. Here, beyond the top 11 combinations in Table 3, the remaining all have posterior probabilities less than $7.0\times 10^{-3}$. Such combinations can thus be safely ignored for debugging, which is indeed the correct decision here. This notion of confidence is not provided by existing methods, which may cause greatly delayed root cause diagnosis, as discussed below.
To contrast, the JMP analysis (Figure 2) yields a much more muddled picture for fault localization. In addition to the top ten tied combinations (which are costly to debug), note that such combinations do not include the true root cause (16), which is a three-factor combination. Thus, even if a test engineer invests resources to debug the top ten combinations, such resources would be wasted and additional resources are needed to identify the true root cause. Furthermore, with this analysis, it is unclear when the tester can confidently stop the debugging of suspicious combinations, as it does not come with a probabilistic quantification of uncertainty. These issues can thus result in costly and delayed fault localization of TCAS; BayesFLo rectifies such issues via a carefully-specified Bayesian modeling framework.
5.2 Case Study 2: JMP Easy DOE Platform
Factor Description Levels
1 Mode Guided, Flexible
2 N_Maximize_Responses 0, 1, 2, 3-5
3 N_Match_Target_Responses 0, 1, 2, 3-5
4 N_Minimize_Responses 0, 1, 2, 3-5
5 N_None_Responses 0, 1, 2, 3-5
6 N_Continuous_Factors 0, 1, 2, 3-5
7 N_Discrete_Numeric_Factors 0, 1, 2, 3-5
8 N_Categorical_Factors 0, 1, 2, 3-5
9 N_Discrete_Numeric_Levels 0, 1, 2, 3-5
10 N_Categorical_Levels 0, 1, 2, 3-5
11 Model_Type Main Effects, Main Effects Interact Corr, Main Effects Interact Uncorr, RSM
12 N_Extra_Runs 0, 2, 3, 4, 5-8
Table 4: Input factors and their corresponding levels for the JMP Easy DOE case study.
Consider next the second case study on the fault localization of JMP’s Easy DOE Platform (JMP Statistical Discovery LLC,, 2023). This platform provides a popular end-to-end framework for experimental design in practical applications, from response and factor definition, model effects and run size specification, design generation, data analysis and results reporting. In testing this platform for deployment, however, the test engineers at JMP observed a failure in its design generation step: the generated design may have fewer runs than the specified run size. To pinpoint the root cause of this failure, the engineers investigated the $I=12$ factors in this module. Table 4 summarizes these factors and their corresponding levels. Here, the first factor controls the usage mode for Easy DOE: “Guided” guides the user through each step of the experimental design process, whereas “Flexible” offers a richer set of options for experienced users. The next four (factors 2 - 5) are response-related factors, which specify the number of responses and their associated goals. The following five (factors 6 - 10) then specify the types and quantities of experiment inputs. Finally, factor 11 indicates the type of statistical model, and factor 12 specifies the number of additional design points in addition to the default sample size from the software.
With this set-up, we construct a strength-2 covering array using the input factors and levels from Table 4, with a run size of $M=36$. Here, for each run, test engineers carefully examined the JMP output against the expected results based on a number of criteria to determine if such outputs align; if not, such a run is marked as a failure. Table 5 shows the corresponding test outcomes. Note that there is only one failure amongst the $M=36$ test runs; such a failure is the aforementioned mismatch in design run size. Using this data, Figure 5 shows the fault localization analysis from the JMP Covering Array module, which we again treat as the state-of-the-art. All eight combinations here are again tied in terms of their suspiciousness; the debugging of all such combinations can thus introduce high costs.
Factor
Run Number 1 2 3 4 5 6 7 8 9 10 11 12 Outcome
1 Guided 0 2 1 0 0 2 2 2 2 Main Effects 3 0
2 Guided 1 3-5 1 2 2 2 1 2 4-6 Main Effects Interact Corr 2 0
3 Guided 3-5 3-5 0 3-5 1 1 1 2 2 Main Effects Interact Corr 5-8 0
4 Flexible 1 0 3-5 3-5 2 2 1 4 4-6 Main Effects 3 0
5 Flexible 0 0 3-5 1 3-5 0 0 0 0 Main Effects Interact Uncorr 0 0
6 Flexible 2 0 3-5 2 1 2 3-5 4 2 RSM 2 0
7 Guided 1 1 0 0 2 1 0 4 0 RSM 0 0
8 Guided 2 0 2 0 3-5 3-5 1 3 4-6 Main Effects Interact Corr 5-8 0
9 Flexible 1 0 0 2 0 2 2 3 4-6 RSM 5-8 0
10 Guided 1 0 1 1 2 3-5 3-5 4 4-6 Main Effects 5-8 0
11 Guided 3-5 1 3-5 2 1 3-5 2 4 4-6 RSM 4 0
12 Flexible 0 3-5 1 3-5 0 3-5 0 4 0 Main Effects 4 0
13 Guided 3-5 2 2 3-5 1 1 0 3 0 Main Effects 3 0
14 Flexible 3-5 1 3-5 0 3-5 0 1 0 2 Main Effects Interact Uncorr 2 1
15 Flexible 0 2 0 3-5 2 2 1 2 4-6 RSM 2 0
16 Flexible 0 1 2 2 0 2 1 3 2 Main Effects 0 0
17 Guided 2 2 0 3-5 1 3-5 3-5 4 3 Main Effects Interact Corr 0 0
18 Guided 1 3-5 3-5 1 1 0 1 0 4-6 RSM 0 0
19 Flexible 1 1 2 1 1 3-5 2 4 2 RSM 0 0
20 Flexible 3-5 0 1 2 1 2 2 3 3 Main Effects 5-8 0
21 Flexible 3-5 1 1 3-5 0 0 3-5 0 4-6 RSM 5-8 0
22 Guided 2 0 3-5 2 1 1 0 2 0 Main Effects 2 0
23 Guided 3-5 2 0 1 2 2 1 3 2 Main Effects 4 0
24 Flexible 0 3-5 3-5 1 0 1 3-5 3 3 Main Effects 4 0
25 Flexible 2 3-5 0 2 1 0 2 0 4-6 Main Effects Interact Corr 3 0
26 Guided 2 1 1 1 3-5 3-5 3-5 2 3 RSM 3 0
27 Guided 0 0 2 0 1 0 1 0 3 Main Effects 4 0
28 Guided 0 1 2 0 0 1 3-5 4 4-6 Main Effects Interact Corr 2 0
29 Guided 1 3-5 1 3-5 3-5 1 2 2 3 Main Effects Interact Corr 4 0
30 Flexible 0 1 2 2 3-5 2 0 2 0 Main Effects 0 0
31 Guided 1 2 2 1 2 3-5 2 3 3 Main Effects 2 0
32 Flexible 1 2 0 2 3-5 0 2 0 2 Main Effects Interact Uncorr 4 0
33 Guided 0 2 3-5 1 3-5 2 0 4 0 Main Effects Interact Corr 5-8 0
34 Guided 2 3-5 2 0 2 0 2 0 0 Main Effects Interact Uncorr 4 0
35 Flexible 2 2 1 3-5 0 1 1 0 2 Main Effects Interact Uncorr 0 0
36 Guided 1 2 2 1 3-5 0 0 0 2 Main Effects Interact Corr 2 0
Table 5: The $M=36$-run CA design and test outcomes for the JMP Easy DOE case study. Here, an outcome of 0 indicates a passed test case, with 1 indicating a failed test case.
Next, we provide a careful prior specification for BayesFLo. Unlike the previous case study, here the test engineers have some prior domain knowledge that can be incorporated. In particular, from the nature of the considered failure (i.e., run size mismatch), the $I=12$ factors from Table 4 can be categorized into three categories of “suspiciousness”. The first category of low suspiciousness includes the four response-based factors (factors 2 - 5), which are least related to the design generation procedure. The second category of high suspiciousness includes the factor N_Extra_Runs, as this specifies the number of additional design runs and is thus most related to the considered failure. Finally, the last category of medium suspiciousness consists of the remaining seven factors. With this, we explore two choices of prior specifications:
  • Prior 1: To reflect these three categories, we set $p_{(i,j)}=0.02$ for factors with low prior suspicion, $p_{(i,j)}=0.04$ for factors with medium suspicion, and $p_{(i,j)}=0.16$ for the factor with high suspicion. This provides an informative prior that captures domain knowledge.
  • Prior 2: This adopts a less informative prior than Prior 1: it only integrates information on the response-based nature of factors 2 - 5. As such $p_{(i,j)}=0.02$ is set for such factors, and $p_{(i,j)}=0.04$ is set for remaining factors. This prior is included to investigate how additional domain knowledge can improve fault localization.
Top Suspicious Combinations (BayesFLo with Prior 1)
Factors Levels Posterior Probability
11, 12 Main Effects Interact Uncorr, 2 0.28
2, 12 3-5, 2 0.14
1, 5 Flexible, 0 0.03
2, 6 3-5, 3-5 0.03
2, 11 3-5, Main Effects Interact Uncorr 0.03
3, 11 1, Main Effects Interact Uncorr 0.03
4, 5 3-5, 0 0.02
2, 5 3-5, 0 0.02
Table 6: Top suspicious combinations from the BayesFLo analysis (with Prior 1) in the JMP Easy DOE case study, ranked by decreasing posterior root cause probabilities. Remaining combinations all have posterior probabilities less than $1.0\times 10^{-2}$. Here, the true root cause is bolded.
Top Suspicious Combinations (BayesFLo with Prior 2)
Factors Levels Posterior Probability
11, 12 Main Effects Interact Uncorr, 2 0.15
2, 12 3-5, 2 0.07
1, 5 Flexible, 0 0.07
2, 5 3-5, 3-5 0.07
2, 11 3-5, Main Effects Interact Uncorr 0.07
3, 11 1, Main Effects Interact Uncorr 0.07
4, 5 3-5, 0 0.04
2, 5 3-5, 0 0.04
Table 7: Top suspicious combinations from the BayesFLo analysis (with Prior 2) in the JMP Easy DOE case study, ranked by decreasing posterior root cause probabilities. Remaining combinations all have posterior probabilities less than $6.0\times 10^{-3}$. Here, the true root cause is bolded.
Figure 5: Top suspicious combinations from the JMP Covering Array analysis for the JMP Easy DOE case study, with its corresponding failure counts from test runs. Here, the true root cause is bolded.
Tables 6 and 7 show, for each prior choice, the top suspicious combinations from BayesFLo ranked by its posterior root cause probabilities. For both prior specifications, the combination with greatest posterior probability is the two-factor combination:
[…]
The probabilistic quantification of suspiciousness from BayesFLo also offers further insights. For the weakly-informative Prior 2, its top combination (which is the true root cause) has a posterior probability of 0.15, followed by seven combinations with posterior probabilities of 0.07 and 0.04; its remaining combinations all have probabilities less than $6.0\times 10^{-3}$. For the more informative Prior 1, its top combination (the true root cause) has a higher posterior probability of 0.28, followed by seven combinations with posterior probabilities of 0.14, 0.03 and 0.02; its remaining combinations all have probabilities less than $1.0\times 10^{-2}$. Thus, despite both priors identifying the true root cause as its most suspicious combination, the integration of more domain knowledge in Prior 1 provides heightened suspiciousness for this combination, which is as desired.
To contrast, the JMP analysis (Figure 5) is likewise more muddled for fault localization. Here, its top eight combinations are all tied, and the debugging of all such combinations to identify the true root cause can be costly. Its lack of probabilistic quantification on combination suspiciousness can similarly lead to a lack of confidence for fault localization. BayesFLo addresses such issues via a flexible probabilistic framework that integrates domain and structural knowledge: it can reduce testing costs by disentangling tied combinations to quickly pinpoint the true root cause for confident fault localization.
6 Conclusion
We proposed a new BayesFLo framework for Bayesian fault localization of complex software systems. Existing fault localization methods largely do not provide a probabilistic quantification of risk on potential root causes nor permit the integration of domain and/or structural knowledge from test engineers, which results many tied suspicious combinations and thus costly fault diagnosis. BayesFLo addresses such limitations via a new Bayesian framework on fault localization, which embeds the combination hierarchy and heredity principles (Lekivetz and Morgan,, 2021) to capture the structured nature of software root causes. One challenge is the computation of posterior root cause probabilities for BayesFLo, which can be infeasible even for small systems. We thus developed a new algorithmic framework for computing the desired posterior probabilities, leveraging recent tools from integer programming and graph representations. We then demonstrate the effectiveness of BayesFLo over the state-of-the-art in two case studies on the Traffic Alert Collision Avoidance System and the JMP Easy DOE platform.
Given promising results, there are many immediate avenues for future work. One direction is the use of the BayesFLo framework for sequential design of subsequent test sets. This adaptive testing of software, facilitated by the proposed Bayesian model, can greatly accelerate the discovery of bugs in complex systems. Another direction is the extension of BayesFLo for fault localization of systems with mixed (i.e., continuous and discrete) factors. Such a setting would be more complex, as it requires the probabilistic modeling of the fault response surface; recent work in Chen et al., (2022) appears to be useful for this goal. Finally, in the rare scenario where one encounters multiple failure types, a reasonable approach may be a Bayesian fault localization framework with categorical responses that can pinpoint root causes for different failure types; we will explore this in future work.
Acknowledgements. This work was supported by NSF CSSI 2004571, NSF DMS 2210729, NSF DMS 2220496, NSF DMS 2316012 and DE-SC0024477.
Appendix A Proof for Proposition 1
Proof: We first define the following two index sets for failed test cases:
$\begin{split}\mathcal{M}_{(\mathbf{i},\mathbf{j})}&=\{m=1,\cdots,M:y_{m}=1,(\mathbf{i},\mathbf{j})\in\mathcal{C}_{{\rm TF},m}\},\\ \mathcal{M}_{-(\mathbf{i},\mathbf{j})}&=\{m=1,\cdots,M:y_{m}=1,(\mathbf{i},\mathbf{j})\notin\mathcal{C}_{{\rm TF},m}\}.\\ \end{split}$
For a given tested-and-failed (TF) combination $(\mathbf{i},\mathbf{j})$, $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ and $\mathcal{M}_{-(\mathbf{i},\mathbf{j})}$ split the failed test cases into cases that either involve or do not involve $(\mathbf{i},\mathbf{j})$.
Next, we define the following events:
$\begin{split}E_{\text{P}}&=\{\text{$Z_{c}=0$ for all $c\in\mathcal{C}_{\rm TP}$}\},\\ E_{(\mathbf{i},\mathbf{j})}&=\{\textup{for each $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m}\setminus\mathcal{C}_{\rm TP}$ such that $Z_{c}=1$}\},\\ E_{-(\mathbf{i},\mathbf{j})}&=\{\textup{for each $m^{\prime}\in\mathcal{M}_{-(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m^{\prime}}\setminus\mathcal{C}_{\rm TP}$ such that $Z_{c}=1$}\}.\end{split}$
Here, $E_{(\mathbf{i},\mathbf{j})}$ is the event that all failures in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ can be explained by some TF combination as a root cause, and $E_{-(\mathbf{i},\mathbf{j})}$ is the event that all failures in $\mathcal{M}_{-(\mathbf{i},\mathbf{j})}$ can be explained by some TF combination as a root cause. Thus, the observed test data $\mathcal{D}$ is equivalent to the intersection of the events $E_{\text{P}}$, $E_{(\mathbf{i},\mathbf{j})}$ and $E_{-(\mathbf{i},\mathbf{j})}$.
As $(\mathbf{i},\mathbf{j})$ is not contained in $E_{\text{P}}$ and $E_{-(\mathbf{i},\mathbf{j})}$ by construction, it follows that $Z_{(\mathbf{i},\mathbf{j})}\perp E_{\text{P}}$ and $Z_{(\mathbf{i},\mathbf{j})}\perp E_{-(\mathbf{i},\mathbf{j})}$. As such, the desired posterior root cause probability can be simplified as:
$\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|\mathcal{D})$ $=\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|E_{(\mathbf{i},\mathbf{j})}).$
Next, note that $\mathbb{P}(E_{(\mathbf{i},\mathbf{j})}|Z_{(\mathbf{i},\mathbf{j})}=1)=1$ as $(\mathbf{i},\mathbf{j})$ is contained in every failure case $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$. We thus get:
$\begin{split}\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1|E_{(\mathbf{i},\mathbf{j})})&=\frac{\mathbb{P}(Z_{(\mathbf{i},\mathbf{j})}=1)\cdot 1}{\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})}\\ &=\frac{p_{(\mathbf{i},\mathbf{j})}}{\mathbb{P}(E_{(\mathbf{i},\mathbf{j})})},\end{split}$
which is as desired. ∎
Appendix B Proof for Proposition 2
Proof: Recall that $E_{(\mathbf{i},\mathbf{j})}$ is defined as:
$E_{(\mathbf{i},\mathbf{j})}=\{\textup{for each $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$, there exists some $c\in\mathcal{C}_{{\rm TF},m}\setminus\mathcal{C}_{{\rm TP}}$ such that $Z_{c}=1$}\}.$
In words, this is the event that each failed test case in $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ can be explained by some TF combination as a root cause. Define $F_{(\mathbf{i},\mathbf{j})}$ as the event:
$F_{(\mathbf{i},\mathbf{j})}=\{\{Z_{c}=1\textup{ for all }c\in\tilde{\mathcal{C}}\}\textup{, for at least one minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$}\}.$
We wish to show that $E_{(\mathbf{i},\mathbf{j})}=F_{(\mathbf{i},\mathbf{j})}$.
Consider first $E_{(\mathbf{i},\mathbf{j})}\subseteq F_{(\mathbf{i},\mathbf{j})}$. This must be true, since if we collect all TF combinations $c$ with $Z_{c}=1$ from $E_{(\mathbf{i},\mathbf{j})}$, they contain at least one minimal cover $\tilde{\mathcal{C}}$, as all failures $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ are covered. This suggests that every arbitrary element in $E_{(\mathbf{i},\mathbf{j})}$ is also an element in $F_{(\mathbf{i},\mathbf{j})}$. Consider next $F_{(\mathbf{i},\mathbf{j})}\subseteq E_{(\mathbf{i},\mathbf{j})}$. This must also be true, since if we take a minimal cover $\tilde{\mathcal{C}}$ of $\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ with $Z_{c}=1$ for all $c\in\tilde{\mathcal{C}}$ from $F_{(\mathbf{i},\mathbf{j})}$, then for every $m\in\mathcal{M}_{(\mathbf{i},\mathbf{j})}$ there exists at least one $c\in\tilde{\mathcal{C}}$ that explains this failure. Thus every arbitrary element in $F_{(\mathbf{i},\mathbf{j})}$ is also an element in $E_{(\mathbf{i},\mathbf{j})}$. This proves the proposition.